# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ 4.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how small LMs and LLMs perform under different hyper-parameter settings. Specifically, we used LoRA to tune a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {16, 32, 64, 640, 1,280, 2,560}. At the same time, we kept other hyper-parameters of LoRA constant, with the LoRA alpha set to 16 and the LoRA dropout set to 0.1. The experiment results are presented in Figure 7.
Figure 7. F1-scores and MCC of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA on detecting CC and CM with different LoRA rank values
The results in Figure 7 show that both CodeBERT and DeepSeek-Coder-1.3B achieve their best performance in detecting CC when fine-tuned with LoRA at an $r$ value of 32. For CM, GraphCodeBERT performs best at a LoRA $r$ value of 32, while DeepSeek-Coder-1.3B reaches its peak performance at $r=64$. As the $r$ value increases from 16 to 2,560, the effectiveness of code smell detection initially increases, then decreases, and finally rises again at $r=2,560$, but it does not exceed the level at $r=16$.
This pattern can be explained by the increase in trainable parameters as $r$ rises, which initially enhances the performance by allowing the models to learn more complex data patterns. However, beyond a certain threshold, the challenge of tuning these additional parameters effectively, especially with relatively limited amount of data, leads to diminishing returns. Without pre-training for these extra parameters, the models are prone to overfitting, where they may learn noise or irrelevant patterns, limiting their generalization capability. Consequently, as the number of trainable parameters grows beyond what the available training data can support, the models’ performance may plateau or even decline. At $r=2,560$, the models appear to utilize the expanded parameter space somewhat more effectively, leading to a slight performance recovery. However, this does not surpass the performance at $r=16$, suggesting that the added parameters introduce more complexity than the data can justify. These results highlight that increasing the number of trainable parameters does not always lead to linear performance gains.
Answer to RQ2: Increasing the number of trainable parameters by adjusting hyper-parameters does not necessarily ensure effectiveness enhancement for code smell detection task. In fact, the effectiveness often declines as the number of trainable parameters continues to increase.
@@ 4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
To simulate the limited availability of data in low-resource scenarios, following previous works (wang2023one, ; liu2023empirical, ), we randomly selected 100, 200, 500, and 1,000 samples from our datasets while maintaining the proportions of the original data distribution. In this RQ, we employed the four PEFT methods and full fine-tuning on a small LM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B) to conduct the experiments. The experiment results are shown in Table 7.
@@
Table 7. Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC and CM in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
Model #Training Samples Method CC CM
F1 MCC F1 MCC
GraphCodeBERT 100 Full FT 44.86% 0.00% 83.87% 69.52%
Prompt Tuning 47.18% 6.84% 80.25% 61.85%
Prefix Tuning 63.94% 32.85% 76.33% 52.67%
LoRA 57.34% 25.28% 82.44% 64.96%
$(IA)^{3}$ 69.63% 39.47% 81.32% 65.84%
200 Full FT 52.48% 21.18% 86.26% 73.93%
Prompt Tuning 67.05% 34.12% 83.26% 67.26%
Prefix Tuning 72.47% 46.00% 81.15% 62.49%
LoRA 66.61% 41.44% 85.20% 71.60%
$(IA)^{3}$ 71.92% 44.12% 84.05% 68.31%
500 Full FT 67.78% 43.34% 86.69% 74.50%
Prompt Tuning 62.44% 27.96% 80.20% 60.40%
Prefix Tuning 69.92% 41.74% 83.26% 66.72%
LoRA 71.08% 45.30% 84.92% 69.84%
$(IA)^{3}$ 67.87% 39.54% 84.25% 69.17%
1,000 Full FT 75.22% 50.81% 85.49% 73.83%
Prompt Tuning 69.63% 39.74% 82.58% 65.30%
Prefix Tuning 73.44% 47.02% 84.16% 68.84%
LoRA 74.76% 49.54% 86.67% 74.16%
$(IA)^{3}$ 63.95% 35.41% 84.07% 68.13%
DeepSeekCoder-1.3B 100 Full FT 44.86% 0.00% 81.48% 68.21%
Prompt Tuning 47.43% 14.34% 78.45% 56.93%
Prefix Tuning 59.82% 20.70% 82.35% 67.39%
LoRA 54.90% 16.98% 83.48% 67.47%
$(IA)^{3}$ 50.93% 8.63% 74.02% 48.52%
200 Full FT 72.76% 50.97% 85.25% 72.17%
Prompt Tuning 60.80% 30.66% 77.63% 55.49%
Prefix Tuning 69.20% 39.65% 82.52% 67.05%
LoRA 66.65% 36.72% 83.18% 66.49%
$(IA)^{3}$ 64.97% 32.56% 77.91% 55.91%
500 Full FT 66.34% 40.34% 83.32% 69.66%
Prompt Tuning 62.13% 30.85% 80.12% 60.25%
Prefix Tuning 71.96% 44.14% 83.91% 68.70%
LoRA 72.94% 45.89% 83.79% 69.70%
$(IA)^{3}$ 63.07% 34.91% 79.17% 58.36%
1,000 Full FT 72.80% 45.63% 82.50% 65.25%
Prompt Tuning 70.24% 42.31% 82.04% 64.10%
Prefix Tuning 70.26% 42.97% 84.43% 69.58%
LoRA 71.84% 44.27% 86.13% 73.00%
$(IA)^{3}$ 73.33% 46.68% 80.33% 60.65%
@@
According to the results reported in Table 7, it is evident that when the data is limited to 100 or 200 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting method-level code smells drops dramatically. This significant decline in performance highlights the challenges small and large LMs face in low-resource scenarios. However, when the number of training samples of CC or CM increases to 1,000, both models show improved performance compared to when only 100 samples are available. This improvement underscores that a larger dataset enables both small and large LMs using PEFT methods to achieve substantially better performance in code smell detection. Moreover, with 1,000 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for both types of method-level code smells. For example, the F1-scores of GraphCodeBERT training with 1,000 samples only degrade by no more than 3.26%, and the MCC decrease by no more than 7.22% compared to training with sufficient data for detecting CM, despite the dataset size being reduced by 90.35% ((10,366 - 1,000) / 10,366). This highlights the benefit of PEFT in resource-limited settings, as it enables the reduction of not only training parameters but also the amount of required training data, while still maintaining strong effectiveness on code smell detection. Notably, when the number of training samples increases to 1,000, GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods demonstrate effectiveness in detecting CC and CM that is on par with, or even exceeds, that of full fine-tuning, further highlighting the superiority of PEFT methods in low-resource scenarios.
Secondly, compared to the results in Table 5 and Table 6, we can find that GraphCodeBERT with full fine-tuning performs better in low-resource scenarios than when there is sufficient data. This improvement can be attributed to the fact that, with ample data, full fine-tuning may lead to overfitting, causing the model to memorize rather than generalize the data. In contrast, in low-resource scenarios, the model is forced to learn more generalizable patterns, thus improving its performance.
@@
Finally, in low-resource scenarios, LoRA tends to be the most effective PEFT method. As indicated by the results of RQ1 (see Section 4.1.4), for both GraphCodeBERT and DeepSeek-Coder-1.3B, fine-tuning with prefix tuning enables them to achieve their best performance compared to the other three PEFT methods. However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest F1-scores or MCC values through LoRA. It suggests that under different resource settings, the most effective PEFT method may vary.
@@
Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 100 or 200 samples. However, as the number of training samples increases, the performance of PEFT methods shows significant improvement, and in some cases, even surpasses that of full fine-tuning.
@@
Prompt Template for Detecting Complex Conditional with GPT-3.5-turbo You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”. Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate. ### Input: [Code Sample] ### Output: [Yes/No]
Prompt Template for Detecting Complex Method with GPT-3.5-turbo You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Method” code smell. If it does, return only “Yes”, else return “No”. Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity. ### Input: [Code Sample] ### Output: [Yes/No]
Traditional heuristic-based tools for code smell detection have been widely used in academia (sharma2021qscored, ; shah2023mining, ; barbosa2020revealing, ), while DL has recently gained widespread adoption in research on code smell detection, achieving impressive performance (zhang2024code, ). In this RQ, we established both heuristic-based (i.e., DesigniteJava) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )), as well as a general-purpose LLM (i.e., GPT-3.5-turobo [^3] https://platform.openai.com/docs/models/o1) which has not been specifically tuned for method-level code smell detection, as baselines. To obtain valid responses from GPT-3.5-turbo in a zero-shot setting (kojima2024large, ), we designed the above prompt templates that include both instructions and definitions of CC or CM to guide ChatGPT in annotating the test sets. GraphCodeBERT with $(IA)^{3}$ for CC and UnixCoder with LoRA for CM are selected for comparison. The results of this evaluation are shown in Table 8.
From the results reported in Table 8, we can find that GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA exhibit significantly better performance than the state-of-the-art code smell detection approaches and GPT-3.5-turbo in detecting CC and CM, respectively. Specifically, for CC, GraphCodeBERT with $(IA)^{3}$ significantly surpasses the effectiveness of the other four baseline methods with an F1-score disparity ranging from 23.41% to 47.94% and an MCC difference ranging from 13.24% to 50.59%. For CM, UnixCoder with LoRA achieves an F1-score improvement ranging from 11.22% to 54.32% and an MCC improvement ranging from 10.28% to 72.91%.
@@
Table 8. Results of GraphCodeBERT with $(IA)^{3}$ and UnixCoder with LoRA compared to state-of-the-art code smell detection approaches on detecting CC and CM (bold: highest values per code smell)
@@
CC DesigniteJava 34.58% 98.74% 51.22% 43.67%
DeepSmells 21.61% 47.17% 29.64% 6.32%
AE-Dense 42.90% 53.82% 47.74% 15.20%
GPT-3.5-turbo 42.86% 73.58% 54.17% 42.66%
GraphCodeBERT with $(IA)^{3}$ 83.33% 74.30% 77.58% 56.91%
CM DesigniteJava 61.88% 99.27% 76.23% 65.92%
DeepSmells 33.82% 56.80% 42.39% 4.65%
AE-Dense 24.33% 55.76% 33.88% 9.78%
GPT-3.5-turbo 81.35% 73.06% 76.98% 67.28%
UnixCoder with LoRA 86.99% 90.66% 88.20% 77.56%
Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods exhibit notable performance improvements on identifying CC and CM, highlighting the effectiveness of PEFT methods for code smell detection.

### v2

@@ 4.2. RQ2: What is the effectiveness of PEFT methods with different hyper-parameters?
4.2.1. Experiment Setup
Given the restrictions of our computational resources and popularity of LoRA, we focused on a single PEFT method, LoRA, to investigate how SLMs and LLMs perform under different hyper-parameter settings. Specifically, we used LoRA to tune a SLM (GraphCodeBERT) and an LLM (DeepSeek-Coder-1.3B), adapting the value of LoRA rank $r$ in the range of {8, 32, 64, 640, 1,280, 2,560}. At the same time, we kept other hyper-parameters of LoRA constant, with the LoRA alpha set to 16 and the LoRA dropout set to 0.1. The experiment results are presented in Figure 6.
Figure 6. MCC results of GraphCodeBERT and DeepSeek-Coder-1.3B using LoRA for code smell detection under different rank settings: (a) CC; (b) CM; (c) FE; (d) DC
4.2.2. Results of RQ2
The results in Figure 6 show that for both GraphCodeBERT and DeepSeek-Coder-1.3B, the optimal $r$ value varies across different code smell types, indicating that the effectiveness of LMs on code smell detection is sensitive to the choice of LoRA configuration. Since the LoRA $r$ value directly controls the number of trainable parameters, this observation implies that different types of code smells may require different levels of model adaptability.
Furthermore, increasing the LoRA $r$ value from 8 to 2,560, which corresponds to an increase in the number of trainable parameters, does not uniformly lead to improved performance. Although a higher $r$ value increases the number of trainable parameters, which provides the model with greater capacity to learn, the MCC scores for detecting the four code smell types do not show a monotonically increasing trend. These findings indicate that expanding the number of trainable parameters alone does not guarantee better effectiveness, emphasizing the importance of careful tuning of the LoRA configuration rather than relying solely on scaling the model.
Answer to RQ2: Increasing the number of trainable parameters by adjusting hyper-parameters does not necessarily ensure effectiveness enhancement for code smell detection task. Additionally, the effectiveness of LMs in detecting different code smells is sensitive to the configuration of PEFT methods (e.g., LoRA).
@@ 4.3. RQ3: What is the effectiveness of PEFT methods in low-resource scenarios?
4.3.1. Experiment Setup
To simulate low-resource scenarios, following previous studies (wang2023one, ; liu2023empirical, ), we randomly selected 50, 100, 250, and 500 samples for CC, CM, and DC, respectively, based on the size of our dataset. Note that each subset is balanced, containing an equal number of positive and negative samples. We excluded FE from this RQ, as the total number of available FE samples in our training set is already below 500 (see Section 3.3.6), making it inherently a low-resource task. The experiment results are shown in Table 8.
@@ 4.3.1. Experiment Setup
Table 8. Results of GrahphCodeBERT and DeepSeek-Coder-1.3B with different PEFT methods and full fine-tuning on detecting CC, CM, and DC in low-resource scenarios (bold: highest values per scenario for a model, underline: highest values among all scenarios for a model)
Model #Training Samples Method CC CM DC
GraphCodeBERT 50 Full FT 19.23% 54.61% 62.76%
Prompt Tuning -0.90% 38.19% 59.51%
Prefix Tuning 6.47% 31.64% 50.79%
LoRA 7.29% 44.99% 62.51%
$(IA)^{3}$ 11.84% 45.15% 61.40%
100 Full FT 24.67% 51.58% 68.28%
Prompt Tuning 3.91% 39.03% 61.49%
Prefix Tuning 15.32% 38.87% 54.19%
LoRA 23.14% 46.98% 68.71%
$(IA)^{3}$ 23.66% 43.99% 65.62%
250 Full FT 28.38% 55.36% 64.93%
Prompt Tuning 13.05% 42.03% 64.69%
Prefix Tuning 20.02% 47.02% 62.19%
LoRA 31.11% 51.91% 66.79%
$(IA)^{3}$ 28.49% 52.45% 64.05%
500 Full FT 34.85% 57.03% 67.64%
Prompt Tuning 18.87% 47.50% 63.41%
Prefix Tuning 25.87% 53.05% 64.63%
LoRA 36.52% 58.31% 67.83%
$(IA)^{3}$ 34.14% 56.22% 68.58%
DeepSeek-Coder-1.3B 50 Full FT 11.38% 42.46% 68.08%
Prompt Tuning 5.88% 25.99% 48.34%
Prefix Tuning -3.70% 43.16% 41.93%
LoRA -1.48% 36.74% 71.84%
$(IA)^{3}$ 1.54% 17.95% 50.29%
100 Full FT 12.27% 24.94% 74.57%
Prompt Tuning 13.05% 33.00% 52.83%
Prefix Tuning 0.68% 45.19% 52.24%
LoRA 10.96% 46.15% 74.01%
$(IA)^{3}$ 6.55% 20.21% 59.53%
250 Full FT 26.69% 48.02% 74.61%
Prompt Tuning 22.18% 38.16% 60.06%
Prefix Tuning 14.87% 45.18% 61.13%
LoRA 20.53% 55.47% 75.33%
$(IA)^{3}$ 15.28% 35.25% 65.16%
500 Full FT 29.33% 58.89% 69.70%
Prompt Tuning 22.09% 40.15% 65.99%
Prefix Tuning 17.72% 46.98% 60.89%
LoRA 32.03% 59.06% 69.57%
$(IA)^{3}$ 9.86% 43.50% 70.78%
4.3.2. Results of RQ3
@@
According to the results reported in Table 8, it is evident that when the data is limited to 50 samples, the effectiveness of both GraphCodeBERT and DeepSeek-Coder-1.3B with PEFT methods in detecting code smells drops, with CC experiencing the most pronounced performance degradation. This significant decline in performance highlights the challenges small and large LMs face in low-resource scenarios. However, when the number of training samples of CC, CM, or DC increases to 500, both models show improved performance compared to when only 50 samples are available. This improvement underscores that a larger dataset enables both small and large LMs using PEFT methods to achieve substantially better performance in code smell detection. Moreover, with 500 samples, GraphCodeBERT and DeepSeekCoder-1.3B with PEFT methods are able to yield reasonably good results compared to training with the original datasets for code smell detection. When the number of training samples reaches 250 or 500, both models with PEFT techniques are able to outperform their fully fine-tuned counterparts.
@@
Additionally, in low-resource scenarios, LoRA tends to be the most effective PEFT method. As indicated by the results of RQ1 (see Section 4.1.6), for GraphCodeBERT, prefix tuning and $(IA)^{3}$ deliver superior performance, while for DeepSeek-Coder-1.3B, $(IA)^{3}$ consistently yields the best performance across different code smells. However, when data is scarce, GraphCodeBERT and DeepSeek-Coder-1.3B often attain their highest MCC values through LoRA. It suggests that under different resource settings, the most effective PEFT method may vary.
@@
Answer to RQ3: In low-resource scenarios, the effectiveness of PEFT methods for code smell detection diminishes considerably when data is limited to 50 samples. However, as the number of training samples increases, the performance of PEFT methods shows significant improvement. Moreover, when the number of training samples reaches 250 or 500, several PEFT techniques are able to outperform full fine-tuning, demonstrating their potential even in moderately low-resource settings.
Table 9. Prompt 1 for GPT-4o-mini and DeepSeek-v3 in detecting CC, CM, FE, and DC
Code Smell Prompt Template*
CC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Conditional” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Complex Conditional”: This code smell occurs when a conditional statement is excessively intricate.
### Input: [Code Sample]
### Output: [Yes/No]
CM You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Complex Method” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Complex Method”: This code smell occurs when a method exhibits high cyclomatic complexity.
### Input: [Code Sample]
### Output: [Yes/No]
FE You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Feature Envy” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Feature Envy”: This code smell occurs when a method is more interested in a class other than the one it actually is in.
### Input: [Code Sample]
### Output: [Yes/No]
DC You are a professional Java programmer. Given a piece of Java code, determine whether it contains “Data Class” code smell. If it does, return only “Yes”, else return “No”.
Definition of “Data Class”: This code smell occurs when a class contains only fields and basic methods for accessing them.
### Input: [Code Sample]
### Output: [Yes/No]
  • * Prompt 1 shown in this table was curated by the authors. Prompt 2 and Prompt 3, which were generated by ChatGPT, are available in our replication package (replicationPackage, ).
@@
4.4.1. Experiment Setup
Figure 7. Results of DeepSeek-v3 and GPT-4o-mini on detecting CC, CM, FE, and DC under ICL
In this RQ, we established both heuristic-based (i.e., DesigniteJava and PMD) and best-in-class DL detectors (i.e., DeepSmells (ho2023fusion, ) and AE-Dense (sharma2021code, )) as baselines for comparison. Additionally, In-Context Learning (ICL) (tom2020language, ) has demonstrated impressive performance in various SE tasks, such as code generation (gao2024what, ; jia2024acecoder, ). To evaluate its potential in code smell detection, we explored the effectiveness of two recent and powerful general-purpose LLMs, GPT-4o-mini (hurst2024gpt4o, ) and DeepSeek-v3 (liu2024deepseek, ), under both zero-shot and few-shot settings. We selected these two models because they represent the latest advancements in commercial and open-weight general-purpose LLMs, respectively, and have demonstrated strong performance across a range of natural language and programming-related tasks (rasheed2024task, ; bruni2025benchmarking, ; liu2024deepseek, ; jahin2025unveiling, ).
Specifically, we first designed a prompt template (i.e., prompt 1) for each of the four types of code smells examined in this study, incorporating both instructional text and smell definitions, as shown in Table 9. We then used ChatGPT (ChatGPT, ) to generate two additional variants for each template (i.e., prompt 2 and prompt 3), resulting in three distinct prompts per code smell type. For the zero-shot setting, these prompts were directly provided to GPT-4o-mini and DeepSeek-v3 without any labeled examples. In the few-shot setting, example pairs were randomly sampled from the training set, where each pair consisted of one code sample with a code smell and one without, along with their corresponding labels. These examples were combined with the base prompts to form few-shot configurations containing one, two, or three example pairs. Due to resource constraints (e.g., limited API usage budget), we restricted the few-shot setting to include no more than three example pairs per prompt. As a result, each code smell type yields a total of 12 unique prompt-example configurations, derived from three prompt variants combined with four levels of contextual information (zero-shot, plus one-, two-, and three-pair few-shot settings).
This ICL setup ensures diversity in both prompt formulations and example composition, allowing for a more comprehensive evaluation of general-purpose LLMs and enabling a fair comparison with our proposed PEFT-based methods in the context of code smell detection. Figure 7 presents the performance of GPT-4o-mini and DeepSeek-v3 on code smell detection in both zero- and few-shot settings. Table 10 summarizes the final results for RQ4, where we report only the best-performing configurations for both GPT-4o-mini and DeepSeek-v3 under ICL across all prompts and example variants.
4.4.2. ICL
It can be observed from Figure 7 that DeepSeek-v3 generally demonstrates better effectiveness in code smell detection compared to GPT-4o-mini. Additionally, both models achieve better performance under few-shot settings than in the zero-shot setting. However, their performance does not consistently improve as the number of example pairs increases.
4.4.3. Comparison with State-of-the-Art Approaches
@@
Table 10. Results of PEFT methods compared to state-of-the-art code smell detection approaches on detecting CC, CM, FE, and DC (bold: highest values per code smell)
@@
CC DesigniteJava 64.01% 76.96% 69.89% 34.41%
DeepSmells 51.06% 100.00% 67.61% 6.63%
AE-Dense 51.13% 49.10% 50.10% 2.92%
GPT-4o-mini 71.48% 64.46% 61.29% 35.24%
DeepSeek-v3 72.38% 71.74% 71.53% 44.12%
StarCoderBase-3B with $(IA)^{3}$ 75.18% 74.78% 74.68% 49.96%
CM DesigniteJava 64.69% 98.61% 78.12% 52.58%
DeepSmells 63.64% 79.87% 70.83% 33.35%
AE-Dense 50.54% 52.05% 51.29% 3.05%
GPT-4o-mini 79.06% 67.25% 63.54% 44.78%
DeepSeek-v3 79.74% 73.78% 72.40% 53.19%
DeepSeek-Coder-6.7B with $(IA)^{3}$ 83.52% 83.36% 83.34% 66.88%
FE DesigniteJava 73.91% 54.84% 62.96% 36.73%
DeepSmells 97.10% 51.54% 67.34% 7.48%
AE-Dense 50.82% 100.00% 67.39% 12.80%
GPT-4o-mini 77.53% 77.42% 77.40% 54.95%
DeepSeek-v3 74.29% 74.19% 74.17% 48.49%
DeepSeek-Coder-1.3B with $(IA)^{3}$ 83.87% 83.87% 83.87% 67.74%
DC PMD 76.29% 89.74% 82.47% 62.83%
DeepSmells 87.72% 67.68% 76.41% 49.55%
AE-Dense 52.79% 94.36% 67.70% 16.18%
GPT-4o-mini 84.87% 82.51% 82.21% 67.35%
DeepSeek-v3 87.57% 87.57% 87.57% 75.15%
GraphCodeBERT with prefix tuning 90.95% 90.03% 89.97% 79.26%
Based on the results presented in Table 10, we can find that LMs fine-tuned with PEFT techniques consistently demonstrate the highest effectiveness among all existing methods for code smell detection. Additionally, among conventional baseline approaches, DeepSeek-v3 achieves the most competitive overall performance. While it shows a significant performance gap compared to LMs tuned with PEFT methods for CM and FE detection (with MCC differences exceeding 10%), its performance in CC and DC detection remains relatively strong, with only a 4%-5% reduction in MCC values, demonstrating its relative strength in these specific code smell types.
Furthermore, our results indicate that traditional DL-based methods (DeepSmells and AE-Dense) exhibit substantially inferior performance compared to both PEFT and the other baseline approaches. This finding underscores the limitations of conventional DL models in code smell detection, particularly when compared to more advanced techniques like PEFT techniques and ICL.
Answer to RQ4: In comparison with existing code smell detection approaches, LMs tuned with PEFT methods reveal notable performance improvements, highlighting the effectiveness of PEFT methods for code smell detection.